Kris mixes plain marshmallows worth 80 cents per gram with flavored marshmallows worth 95 cents per gram to make 120 grams of marshmallows to be sold at 85 cents a gram. How many grams of flavored marshmallows does she use?
Answer by lwsshak3(11628)

Kris mixes plain marshmallows worth 80 cents per gram with flavored marshmallows worth 95 cents per gram to make 120 grams of marshmallows to be sold at 85 cents a gram. How many grams of flavored marshmallows does she use?
***
let x=amount of flavored marshmallows used
120-x=amount of plain marshmallows used
..
.95x+.80(120-x)=.85*120
.95x+96-.80x=102
.15x=102-96=6
x=40
amount of flavored marshmallows used=40 grams